OpenStudy (anonymous):

If events A and B are independent, find P(A|B) a. P(A) b. 1 c. 0 d. P(A) / P(B)

myininaya (myininaya):

\[P(A n B) =P(A) \cdot P(B) \text{ if A and B are independent }\]

myininaya (myininaya):

\[P(A|B)=\frac{P(A n B)}{P(B)}\]

myininaya (myininaya):

put what i gave you and you will find what you seek

OpenStudy (anonymous):

So it will be P(A) correct?

OpenStudy (anonymous):

@myininaya

myininaya (myininaya):

yep